# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ F.5 Ablation study: integrating wavelet filters to HGCNs without iterative learning scheme
It is important to note that Haar wavelet filters are induced by a tree structure [Gavish et al., 2010]. However, in the case of hierarchical graph data, the structure is not always a tree. To quantify how closely a graph resembles a tree, one can use the notion of $\delta$-hyperbolicity [Gromov, 1987]. A tree has $\delta=0$, whereas hierarchical graphs that deviate from tree-like geometry exhibit larger $\delta$-hyperbolicity values, indicating lower tree-likeness or increased curvature. For the hierarchical graph datasets considered in our experiments, following prior work on HGCNs [Chami et al., 2019, Dai et al., 2021], the measured $\delta$-hyperbolicity values are as follows: DISEASE ($\delta=0$), AIRPORT ($\delta=1$), PUBMED ($\delta=3.5$), and CORA ($\delta=11$). Among them, only DISEASE forms a tree, allowing the Haar wavelet to be applied directly. For the other datasets, which are not trees, we can either convert the graph into a tree before applying Haar wavelets or construct Haar-like wavelets using the multiscale tree approximation framework described in [Gavish et al., 2010].
Table 9: Performance comparison of HGCNs integrated with wavelet filtering variants (“-MST”, “-HC”, and “-Wavelet”) versus our full method that combines iterative learning with wavelet filtering. Results are reported for link prediction and node classification tasks across hierarchical graph datasets. ROC AUC for LP, F1 score for the DISEASE dataset, and accuracy for the AIRPORT, PUBMED, and CORA datasets for NC tasks. Arrows ()($\uparrow$)( ↑ ) denote cases with improvement. The highest performance is marked in bold, and the second-highest is underlined.
@@
HGCNs [Chami et al., 2019] 90.8$\pm$0.3 74.5$\pm$0.9 96.4$\pm$0.1 90.6$\pm$0.2 96.3$\pm$0.0 80.3$\pm$0.3 92.9$\pm$0.1 79.9$\pm$0.2
HGCNs-MST - - 92.3$\pm$0.2 88.4$\pm$0.4 91.7$\pm$0.2 77.9$\pm$0.3 90.4$\pm$0.3 77.6$\pm$0.6
HGCNs-HC - - 90.9$\pm$0.2 87.9$\pm$0.3 92.6$\pm$0.3 76.6$\pm$0.4 91.3$\pm$0.5 76.4$\pm$0.7
HGCNs-Wavelet 91.2$\pm$0.6 ($\uparrow$) 76.7$\pm$1.0 ($\uparrow$) 96.0$\pm$0.2 90.6$\pm$0.3 96.4$\pm$0.1 ($\uparrow$) 80.1$\pm$0.4 93.0$\pm$0.2 ($\uparrow$) 80.4$\pm$0.3 ($\uparrow$)
H2H-GCN [Dai et al., 2021] 97.0$\pm$0.3 88.6$\pm$1.7 96.4$\pm$0.1 89.3$\pm$0.5 96.9$\pm$0.0 79.9$\pm$0.5 95.0$\pm$0.0 82.8$\pm$0.4
H2H-GCN-MST - - 90.5$\pm$0.3 88.6$\pm$0.4 91.4$\pm$0.7 78.6$\pm$0.6 92.4$\pm$0.3 81.3$\pm$0.5
H2H-GCN-HC - - 92.1$\pm$0.2 88.9$\pm$0.4 92.6$\pm$0.4 79.8$\pm$0.2 92.5$\pm$0.5 81.7$\pm$0.4
H2H-GCN-Wavelet 97.2$\pm$0.5 ($\uparrow$) 87.9$\pm$1.8 96.4$\pm$0.4 90.1$\pm$0.4 ($\uparrow$) 95.2$\pm$0.1 80.2$\pm$0.5 ($\uparrow$) 94.7$\pm$0.2 82.4$\pm$0.3
HGCN-Alg. 1 93.2$\pm$0.6 ($\uparrow$) 87.9$\pm$0.7 ($\uparrow$) 93.7$\pm$0.2 ($\uparrow$) 89.9$\pm$0.4 94.1$\pm$0.7 81.7$\pm$0.2 ($\uparrow$) 93.1$\pm$0.1 ($\uparrow$) 82.9$\pm$0.3 ($\uparrow$)
HGCN-Alg. 2 98.4$\pm$0.4 ($\uparrow$) 89.4$\pm$0.3 ($\uparrow$) 97.2$\pm$0.1 ($\uparrow$) 92.1$\pm$0.3 ($\uparrow$) 97.2$\pm$0.2 ($\uparrow$) 83.6$\pm$0.4 ($\uparrow$) 96.9$\pm$0.3 ($\uparrow$) 83.9$\pm$0.2 ($\uparrow$)
For the naive transformation, we consider two strategies to approximate the graph with a tree structure: (i) we construct an MST by connecting all nodes using edges with the smallest cumulative weights; (ii) we apply agglomerative clustering to iteratively merge nodes and form a hierarchical tree, as in standard hierarchical clustering. These transformations alter the original graph topology, but the resulting tree is intended to closely approximate the structure of the original graph. Once the tree is constructed, it can be used to define a Haar wavelet basis and build the corresponding wavelet filters, as described in Sec. 4.2, which are then applied to the node attributes (i.e., features). To construct Haar-like wavelets using the multiscale tree approximation framework of [Gavish et al., 2010], we first build a hierarchical partition tree over the data using a suitable distance metric (e.g., shortest path distance). Each wavelet basis function is then defined over a node in the tree by assigning opposite-signed, normalized values to its children. We denote wavelet filtering applied to the node attributes (i.e., features) using these approximations, without our iterative learning scheme, by “-MST” (for MST-based tree), “-HC” (for hierarchical clustering), and “-Wavelet” (for multiscale tree approximation). We incorporate these variants into HGCNs [Chami et al., 2019, Dai et al., 2021] and evaluate their performance on link prediction and node classification tasks.
@@
Table 10: Comparison of classification accuracy across word-document and scRNA-seq datasets for three variants: Alg. 1, Alg. 1-Wavelet (post-hoc wavelet filtering after convergence), and Alg. 2 (wavelet filtering applied iteratively during learning). Arrows ()($\uparrow$)( ↑ ) indicate improvements. We see that while post-hoc filtering offers minor improvements, iterative integration yields consistently better performance.
@@ F.9 Runtime analysis
Figure 10: Runtime comparison of our methods and competing baselines.
@@ F.10 Effect of Sinkhorn regularization in the iterative learning scheme
Our method adopts a snowflake-based regularization function, as proposed in [Mishne et al., 2019], which promotes smoothness across both rows and columns of the data matrix. T his regularizer increases monotonically over $[0,\infty)$ and penalizes small differences more strongly, encouraging local consistency. While effective, other regularization strategies may also be considered. One notable alternative is entropic regularization [Cuturi, 2013], which solves the OT problem using matrix scaling.
To explore this, we implement the entropic regularization using the Sinkhorn solver from [Flamary et al., 2021] within our iterative learning framework. We denote these variants as Alg. 1-Sinkhorn and Alg. 2-Sinkhorn. Tab. 11 reports the classification results on scRNA-seq and word-document datasets using the Sinkhorn regularization. In all cases, incorporating the Sinkhorn penalty leads to noticeably lower performance compared to the original versions of our algorithms. The classification accuracy degrades, indicating that the entropic regularization may overly smooth or distort the learned hierarchical structures. This suggests that the snowflake regularization is more suitable in our setting, where the goal is to represent fine-grained hierarchical relationships across both samples and features.
@@ F.11 Co-Clustering performance on gene expression data
Table 12: Clustering performance on gene expression datasets (Breast Cancer and Leukemia) evaluated using Clustering Accuracy (CA), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI). Our methods (Alg. 1 and Alg. 2) outperform all baselines. Alg. 2, which incorporates wavelet filtering, achieves the highest overall performance. Results for BCOT, CCOT, CCOT-GW, and COOT are reported from [Fettal et al., 2022]. OOM indicates methods that ran out of memory.
@@
QUE 71.9$\pm$1.2 45.3$\pm$0.7 39.8$\pm$0.4 72.3$\pm$0.5 56.6$\pm$0.6 42.9$\pm$0.4
WSV 72.4$\pm$3.6 33.8$\pm$2.2 39.0$\pm$0.8 64.3$\pm$4.1 61.7$\pm$2.3 44.8$\pm$5.1
Tree-WSV 73.1$\pm$1.8 36.7$\pm$1.7 41.5$\pm$0.9 66.2$\pm$3.0 60.4$\pm$1.8 42.6$\pm$3.1
BCOT 76.9$\pm$0.0 37.2$\pm$0.0 26.7$\pm$0.0 71.2$\pm$5.4 59.6$\pm$6.9 39.9$\pm$6.3
$\text{BCOT}_{\lambda}$ 84.6$\pm$0.0 48.3$\pm$0.0 46.0$\pm$0.0 80.9$\pm$3.8 70.9$\pm$4.1 55.3$\pm$3.3
CCOT OOM OOM OOM 40.6$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
@@
COOT 63.1$\pm$5.2 5.4$\pm$8.7 -1.2$\pm$2.9 36.2$\pm$2.7 14.0$\pm$3.6 5.4$\pm$3.2
$\text{COOT}_{\lambda}$ 61.5$\pm$0.0 5.4$\pm$0.0 2.2$\pm$0.0 32.5$\pm$3.3 8.7$\pm$2.7 -0.5$\pm$2.1
Alg. 1 85.6$\pm$0.7 49.2$\pm$0.5 50.4$\pm$0.9 75.8$\pm$0.9 64.2$\pm$1.0 54.2$\pm$1.1
Alg. 2 87.2$\pm$0.4 51.3$\pm$0.3 57.6$\pm$0.7 81.2$\pm$1.1 75.4$\pm$0.8 59.7$\pm$1.2
We further evaluate the effectiveness of our methods in clustering tasks using two gene expression datasets from [Feltes et al., 2019]. The first is a breast cancer (BC) dataset containing 42,945 gene expression values across 26 samples divided into two classes. The second is a leukemia (LEU) dataset with 22,283 genes measured across 64 patients divided into five classes. We apply Alg. 1 and Alg. 2 to the data and perform $k$-means clustering [Hartigan and Wong, 1979] on the resulting sample TWD matrices to obtain clustering assignments. We evaluate the OT-based co-clustering baselines including: QUE [Ankenman, 2014], WSV [Huizing et al., 2022], Tree-WSV [Düsterwald et al., 2025], biclustering with optimal transport (BCOT and $\text{BCOT}_{\lambda}$) [Fettal et al., 2022], Co-clustering through OT (CCOT and CCOT-GW) [Laclau et al., 2017], and Co-optimal transport (COOT and $\text{COOT}_{\lambda}$) [Titouan et al., 2020]. We follow the evaluation setup in [Fettal et al., 2022], reporting Clustering Accuracy (CA), Normalized Mutual Information (NMI) [Cai et al., 2008], and Adjusted Rand Index (ARI) [Hubert and Arabie, 1985]. The results of BCOT, CCOT, CCOT-GW, and COOT are reported from [Fettal et al., 2022].
@@
F.12 Incorporating fixed hierarchical distances in iterative learning scheme for HGCNs
Algorithm 4
@@ F.12 Incorporating fixed hierarchical distances in iterative learning scheme for HGCNs
  $l\leftarrow 0$, ${\mathbf{X}}^{(0)}\leftarrow{\mathbf{X}}$, ${\mathbf{Z}}^{(0)}\leftarrow\mathbf{X}^{\top}$ and $\widetilde{\mathbf{W}}_{c}^{(0)}\leftarrow\mathbf{M}_{c}$
@@
   ${\mathbf{X}}^{(l+1)}\leftarrow\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widetilde{\mathbf{W}}_{c}^{(l)}))$ and ${\mathbf{Z}}^{(l+1)}\leftarrow\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\mathbf{M}_{r}))$
   $\widetilde{\mathbf{X}}_{r}^{(l+1)}\leftarrow\left\{\mathbf{r}_{i}^{(l+1)}=\left({\mathbf{X}}_{i,:}^{(l+1)}\right)^{\top}\middle/\left\lVert{\mathbf{X}}_{i,:}^{(l+1)}\right\rVert_{1}\right\}$
   $\widetilde{\mathbf{X}}_{c}^{(l+1)}\leftarrow\left\{\mathbf{c}_{j}^{(l+1)}=\left({\mathbf{Z}}_{j,:}^{(l+1)}\right)^{\top}\middle/\left\lVert{\mathbf{Z}}_{j,:}^{(l+1)}\right\rVert_{1}\right\}$
   $\widetilde{\mathbf{W}}_{r}^{(l+1)}\leftarrow\Phi(\widetilde{\mathbf{X}}_{r}^{(l+1)};\mathcal{T}(\widetilde{\mathbf{W}}_{c}^{(l)}))$ and $\widetilde{\mathbf{W}}_{c}^{(l+1)}\leftarrow\Phi(\widetilde{\mathbf{X}}_{c}^{(l+1)};\mathcal{T}(\mathbf{M}_{r}))$
@@
HGCN-Alg. 1 93.2$\pm$0.6 87.9$\pm$0.7 93.7$\pm$0.2 89.9$\pm$0.4 94.1$\pm$0.7 81.7$\pm$0.2 93.1$\pm$0.1 82.9$\pm$0.3
HGCN-Alg. 4 95.1$\pm$0.5 88.2$\pm$0.9 95.9$\pm$0.1 90.7$\pm$0.3 95.7$\pm$0.4 82.0$\pm$0.3 95.3$\pm$0.1 83.2$\pm$0.4
HGCN-Alg. 2 98.4$\pm$0.4 89.4$\pm$0.3 97.2$\pm$0.1 92.1$\pm$0.3 97.2$\pm$0.2 83.6$\pm$0.4 96.9$\pm$0.3 83.9$\pm$0.2

### v3

@@ F.5 Ablation study: integrating wavelet filters to HGCNs without iterative learning scheme
It is important to note that Haar wavelet filters are induced by a tree structure [43]. However, in the case of hierarchical graph data, the structure is not always a tree. To quantify how closely a graph resembles a tree, one can use the notion of $\delta$-hyperbolicity [107]. A tree has $\delta=0$, whereas hierarchical graphs that deviate from tree-like geometry exhibit larger $\delta$-hyperbolicity values, indicating lower tree-likeness or increased curvature. For the hierarchical graph datasets considered in our experiments, following prior work on HGCNs [21, 45], the measured $\delta$-hyperbolicity values are as follows: DISEASE ($\delta=0$), AIRPORT ($\delta=1$), PUBMED ($\delta=3.5$), and CORA ($\delta=11$). Among them, only DISEASE forms a tree, allowing the Haar wavelet to be applied directly. For the other datasets, which are not trees, we can either convert the graph into a tree before applying Haar wavelets or construct Haar-like wavelets using the multiscale tree approximation framework [43].
For the naive transformation, we consider two strategies to approximate the graph with a tree structure: (i) we construct an MST by connecting all nodes using edges with the smallest cumulative weights; (ii) we apply agglomerative clustering to iteratively merge nodes and form a hierarchical tree, as in standard hierarchical clustering. These transformations alter the original graph topology, but the resulting tree is intended to closely approximate the structure of the original graph. Once the tree is constructed, it can be used to define a Haar wavelet basis and build the corresponding wavelet filters, as described in Sec. 4.2, which are then applied to the node attributes (i.e., features). To construct Haar-like wavelets using the multiscale tree approximation framework [43], we first build a hierarchical partition tree over the data using a suitable distance metric (e.g., shortest path distance). Each wavelet basis function is then defined over a node in the tree by assigning opposite-signed, normalized values to its children. We denote wavelet filtering applied to the node attributes (i.e., features) using these approximations, without our iterative learning scheme, by “-MST” (for MST-based tree), “-HC” (for hierarchical clustering), and “-Wavelet” (for multiscale tree approximation). We incorporate these variants into HGCNs [21, 45] and evaluate their performance on link prediction and node classification tasks.
Table 9: Performance comparison of HGCNs integrated with wavelet filtering variants (“-MST”, “-HC”, and “-Wavelet”) versus our full method that combines iterative learning with wavelet filtering. Results are reported for link prediction and node classification tasks across hierarchical graph datasets. ROC AUC for LP, F1 score for the DISEASE dataset, and accuracy for the AIRPORT, PUBMED, and CORA datasets for NC tasks. Arrows $(\uparrow)$ denote cases with improvement. The highest performance is marked in bold, and the second-highest is underlined.
@@
[…]
@@
Table 10: Comparison of classification accuracy across word-document and scRNA-seq datasets for three variants: Alg. 1, Alg. 1-Wavelet (post-hoc wavelet filtering after convergence), and Alg. 2 (wavelet filtering applied iteratively during learning). Arrows $(\uparrow)$ indicate improvements. We see that while post-hoc filtering offers minor improvements, iterative integration yields consistently better performance.
@@ F.9 Runtime analysis
Figure 10: Runtime comparison of our methods and competing baselines.
@@ F.10 Effect of Sinkhorn regularization in the iterative learning scheme
Our method adopts a snowflake-based regularization function [57], which promotes smoothness across both rows and columns of the data matrix. T his regularizer increases monotonically over $0,\infty)$ and penalizes small differences more strongly, encouraging local consistency. While effective, other regularization strategies may also be considered. One notable alternative is entropic regularization [119], which solves the OT problem using matrix scaling.
To explore this, we implement the entropic regularization using the Sinkhorn solver [155] within our iterative learning framework. We denote these variants as Alg. 1-Sinkhorn and Alg. 2-Sinkhorn. Tab. 11 reports the classification results on scRNA-seq and word-document datasets using the Sinkhorn regularization. In all cases, incorporating the Sinkhorn penalty leads to noticeably lower performance compared to the original versions of our algorithms. The classification accuracy degrades, indicating that the entropic regularization may overly smooth or distort the learned hierarchical structures. This suggests that the snowflake regularization is more suitable in our setting, where the goal is to represent fine-grained hierarchical relationships across both samples and features.
@@ F.11 Co-Clustering performance on gene expression data
Table 12: Clustering performance on gene expression datasets (Breast Cancer and Leukemia) evaluated using Clustering Accuracy (CA), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI). Our methods (Alg. 1 and Alg. 2) outperform all baselines. Alg. 2, which incorporates wavelet filtering, achieves the highest overall performance. Results for BCOT, CCOT, CCOT-GW, and COOT are reported from the work of Fettal et al. [156]. OOM indicates methods that ran out of memory.
@@
[…]
@@
[…]
We further evaluate the effectiveness of our methods in clustering tasks using two gene expression datasets [157]. The first is a breast cancer (BC) dataset containing 42,945 gene expression values across 26 samples divided into two classes. The second is a leukemia (LEU) dataset with 22,283 genes measured across 64 patients divided into five classes. We apply Alg. 1 and Alg. 2 to the data and perform $k$-means clustering [158] on the resulting sample TWD matrices to obtain clustering assignments. We evaluate the OT-based co-clustering baselines including: QUE [52], WSV [59], Tree-WSV [60], biclustering with optimal transport (BCOT and $\text{BCOT}_{\lambda}$) [156], Co-clustering through OT (CCOT and CCOT-GW) [159], and Co-optimal transport (COOT and $\text{COOT}_{\lambda}$) [160]. We follow the evaluation setup [156], reporting Clustering Accuracy (CA), Normalized Mutual Information (NMI) [161], and Adjusted Rand Index (ARI) [162]. The results of BCOT, CCOT, CCOT-GW, and COOT are reported from the work of Fettal et al. [156].
@@
F.12 Incorporating fixed hierarchical distances in the iterative learning scheme for HGCNs
Algorithm 4 Incorporating Fixed Hierarchical Distance for One Mode in Iterative Learning
@@ F.12 Incorporating fixed hierarchical distances in the iterative learning scheme for HGCNs
$l\leftarrow 0$, ${\mathbf{X}}^{(0)}\leftarrow{\mathbf{X}}$, ${\mathbf{Z}}^{(0)}\leftarrow\mathbf{X}^{\top}$, $\widetilde{\mathbf{W}}_{c}^{(0)}\leftarrow\mathbf{M}_{c}$ $\triangleright$ Initialization
@@
${\mathbf{X}}^{(l+1)}\leftarrow\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widetilde{\mathbf{W}}_{c}^{(l)}))$ $\triangleright$ Tree haar wavelet filtering with updated tree
${\mathbf{Z}}^{(l+1)}\leftarrow\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\mathbf{M}_{r}))$ $\triangleright$ Tree haar wavelet filtering with fixed tree
$\widetilde{\mathbf{X}}_{r}^{(l+1)}\leftarrow\left\{\mathbf{r}_{i}^{(l+1)}=\left({\mathbf{X}}_{i,:}^{(l+1)}\right)^{\top}/\|{\mathbf{X}}_{i,:}^{(l+1)}\|_{1}\right\}$
$\widetilde{\mathbf{X}}_{c}^{(l+1)}\leftarrow\left\{\mathbf{c}_{j}^{(l+1)}=\left({\mathbf{Z}}_{j,:}^{(l+1)}\right)^{\top}/\|{\mathbf{Z}}_{j,:}^{(l+1)}\|_{1}\right\}$
$\widetilde{\mathbf{W}}_{r}^{(l+1)}\leftarrow\Phi(\widetilde{\mathbf{X}}_{r}^{(l+1)};\mathcal{T}(\widetilde{\mathbf{W}}_{c}^{(l)}))$ $\triangleright$ Iterative update with updated tree
$\widetilde{\mathbf{W}}_{c}^{(l+1)}\leftarrow\Phi(\widetilde{\mathbf{X}}_{c}^{(l+1)};\mathcal{T}(\mathbf{M}_{r}))$ $\triangleright$ Iterative update with fixed tree
@@
F.13 Evaluating the preprocessing step with additional backbones
To evaluate the effectiveness of our method as a preprocessing step across different neural network backbones, we extend our experiments to include GCN and GraphSAGE, in addition to the hyperbolic models considered in Tab. 3. Tab. 14 reports link prediction and node classification results, where models initialized with our learned hierarchical representations (denoted “-Alg. 2”) consistently outperform their baselines. Improvements are indicated with an upward arrow ($\uparrow$). These findings demonstrate the broad applicability of our approach across diverse neural architectures. Among all models, HGCN-Alg. 2 achieves the highest overall performance, which we attribute to the advantages of hyperbolic representations in modeling hierarchical structures, distinguishing HGCNs from their Euclidean counterparts.
Table 14: Link prediction and node classification results for GCN, GraphSAGE, and HGCNs, with and without initialization using our learned hierarchical representations (“-Alg. 2”).
@@ F.13 Evaluating the preprocessing step with additional backbones
[…]
Table 15: Comparison of representation quality for Independent MST, Co-MST-TWD, and our full method (Alg. 2), measured by the $L_{1}$ norm of Haar coefficients (lower is better). Results are reported as $\mathrm{samples}$ / $\mathrm{features}$ for each dataset. Lower values indicate that the learned hierarchical representations more effectively capture the hierarchical information across the two data modes.
BBCSPORT TWITTER CLASSIC AMAZON ZEISEL CBMC
Independent MST 37.9 / 42.8 82.4 / 31.1 83.6 / 131.4 109.6 / 144.1 194.3 / 281.5 1928.7 / 97.6
Alg. 2 10.1 / 4.8 22.4 / 3.4 37.2 / 19.4 46.6 / 26.6 50.9 / 93.7 489.4 / 45.6
Figure 11: Stepwise hierarchical refinement on the toy user–video recommendation in App. F.1.
F.14 Discussion and comparison with other tree distance metrics
In our framework, TWD plays a central role due to its ability to compare probability distributions supported on hierarchical structures. This is essential for our alternating, joint-learning process, where the hierarchical structure from one mode (e.g., samples) is used to inform and refine the hierarchy in the other mode (e.g., features), and vice versa. That is, the tree structure provides the foundation for expressing hierarchy, while the Wasserstein metric provides the mechanism for distributional comparison and transfer from one mode to the other. This cross-mode interaction is central to our method and relies on the Wasserstein nature of TWD.
TWD differs from other tree distance metrics such as minimum spanning tree (MST) [163], neighbor joining, or low-stretch trees, which typically operate on tree structures or discrete node sets but do not directly support distribution-based comparison needed for cross-mode refinement. To clarify this difference, we consider two alternative setups using MST: (1) using independently constructed MSTs for samples and features without cross-mode interaction, and (2) using an MST as the tree structure in the TWD framework for our iterative learning process (Co-MST-TWD). We compare these alternatives to our full method (Alg. 2) by measuring representation quality using the $L_{1}$ norm of Haar coefficients. Results are shown in Tab. 15 (format: $\mathrm{samples}$ / $\mathrm{features}$). We see that both alternatives yield substantially higher $L_{1}$ norms. These results suggest that the learned joint hierarchical representation from our iterative process better reflects the hierarchical information of the two data modes. We attribute this advantage to the specific design of our proposed method, where the TWD supports iterative cross-mode refinement, which is not feasible with independent tree metrics.
Replacing TWD with other graph-based distances is conceptually possible, but there are important considerations:
  • Most graph-based distances are designed to compare node positions or global connectivity in general graphs, not to measure how distributions are supported on a tree or reveal explicit nested groupings.
  • Gromov-Wasserstein distance [164] could, in principle, be integrated into an iterative scheme as it compares distributions across different metric spaces or graphs. However, such approaches are much more computationally intensive.
  • Another possible direction is to construct a graph from high-dimensional data and transform this general graph into unrolling trees [165]. Further investigation would be needed to determine if and how such transformations can be integrated within our iterative learning process.
F.15 Stepwise hierarchical visualizations in recommendation toy problem
Fig. 11 presents the stepwise visualizations based on the toy video recommendation example in App. F.1 for hierarchical refinement. The iterative refinement process starts with an initial tree structure, which is unstructured since the preliminary grouping is based on initial data relationships ($\mathbf{M}_{c}$ or $\mathbf{M}_{r}$). At this point, major clusters are not well formed, and samples or features from different categories are mixed. After the first iteration, the algorithm uses the current tree from one mode (e.g., users) to compute a TWD for the other mode (e.g., videos). This leads to a newly induced tree that starts to separate the major categories, and major branches start to reflect groupings, but subcategory assignments are not fully sorted. As the algorithm proceeds through subsequent iterations, it alternately refines the trees for users and videos. The updated tree for one mode informs a more accurate computation of TWD in the other mode, enabling finer divisions within the existing branches. Subcategories like action, drama, and sci-fi (for fiction videos) become more distinct and better grouped. After several iterations, the subcategories become more cohesive and consistently organized, and the tree structures stabilize. It illustrates how the trees evolve with each iteration and points out key groupings formed in the learned trees.